OpenStudy (anonymous):

How to find the exact value of cos(cos-1 pi) using calculator??? I keep getting error.

OpenStudy (phi):

Are you sure the problem isn't \[ \cos^{-1}( \cos( \pi)) \] the cosine of an angle gives you a number between -1 and +1 the inverse cosine of a number (if that number is between -1 and +1) gives you an angle. notice that pi is bigger than +1, and the inverse cosine has no (real) answer. In other words, there is no angle where the cosine(angle) = 3.14 (all angles will have a cosine between -1 and +1)

OpenStudy (anonymous):

So in that case would it be 'Not defined'? My teacher gave me this problem maybe it was her mistake.

OpenStudy (phi):

If the problem is \[ \cos^{-1}( \cos( \pi)) \] you don't need a calculator. the cosine of pi is a number (-1, but we don't need to know that). then the inverse cosine of that number returns back pi if the problem is as stated, there is no real solution.

OpenStudy (anonymous):

Thank you!